The entropy of vapourisation of benzene is 85JK-1mol-1.When 117g benzene vaporize at its normal boiling point, the entropy change of surrounding is (assuming reversible process):
(a)-85JK-1
(b)-85×1.5 JK-1
(c)85 ×1.5JK-1
(d) none of these
1
Expert's answer
2018-04-09T03:18:11-0400
∆S=∆S_mol ν=∆S_mol m/M=85 117/78.11=85(1.5) J/K